OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? -2x+5=-12 -6x+__y=__

OpenStudy (alexandervonhumboldt2):

well first lets solve for x: -2x+5=-12 -2x=-17 2x=17 x=17/2=8.5

OpenStudy (pawanyadav):

Let equations a1x+b1y=c1 And. a2x+b2y=c2 Then for coincident a1/a2a2=b1/b2=c1/c2

OpenStudy (pawanyadav):

BTW in question equation should be -2x+5y=-12

OpenStudy (mrnood):

@Pawanyadav Your numbers in your first answer are not correct

OpenStudy (mrnood):

and maybe you should let the asker work out the correct numbers herself

OpenStudy (pawanyadav):

@MrNood